OpenStudy (anonymous):

A piece of solid CO2 with a mass of 7.8 g is in an empty, 4.0L container at 27 C. What is the vapor pressure in the container once the CO2 vaporizes?

OpenStudy (anonymous):

I know I need to use \[P=n(R)(300 K)/4.0L\] and I have a value for R, but I need a value n where n is the number of moles of CO2 in 7.8 g of solid CO2.

OpenStudy (anonymous):

I calculated n to be 0.233, so got a value of 1.42 atm of CO2, but the answer sheet tells me it should be 1.1 atm. I'm making a mistake somewhere... :/

OpenStudy (anonymous):

i got also 1,1 atm, so lets write it down: n= m/M = 7,8 g / 44,009 gmol = 0,177236 mol pV=nRT p= nRT/V p= 0,177236 mol * 8,314 J K^-1mol^-1 * (273 + 27)K / 4*10^-3 m^3 = 110515,8 Pa or 1,1 atm

OpenStudy (anonymous):

Ah, thank you Kryten. I'd made two mistakes. One, I had made a math error in calculating n, and two I used a wrong value for R. Thank you again.

OpenStudy (anonymous):

not a problem ;)